OpenStudy (anonymous):

how do i simplify (y^-1+x^-1)/(xy)^-1

OpenStudy (displayerror):

I'm assuming your equation is this: \[\frac{y^{-1} + x^{-1}}{(xy)^{-1}}\] Remember this rule of exponents: \[a^{-b} = \frac{1}{a^b}\] Applying that rule to this problem: \[\frac{\frac{1}{y} + \frac{1}{x}}{\frac{1}{xy}}\] Recall the use of reciprocals when dividing by fractions: \[(\frac{1}{x} + \frac{1}{y}) * xy\] Distribute the xy term: \[\frac{xy}{x} + \frac{xy}{y}\] Now we can cancel to get the answer: \[y+x\]